Could someone please help me with this one;
10,000 = (SQRT x/360) * 0.15 * 200,000

Found 2 solutions by mananth, unlockmath:
Answer by mananth(16949)

10,000 = (SQRT x/360) * 0.15 * 200,000
10000 /(0.15*200,000) = sqrt(x/360)
1/(0.15* 20)= sqrt(x/360)
1/3=sqrt(x/360)
take square on bothe sides
1/9 =x/360
x=360/9
x=40

Answer by unlockmath(1688)

Hello,
Let's square both sides to get rid of the sqrt so it will look like:
10,000^2 = x/360 * 0.15^2 * 200,000^2 Do the calculations and multiply by 360:
10,000^2(360) =x*.0225*200,000^2 Now Divide both sides by .0225*200,000 gives us:
10,000^2(360)/(.0225*200,000)=x This comes to:
x=40
Since these are huge numbers, it's best to use a scientific calculator.
RJ
WWW.math-unlock.com